### v2

# Efficient Feature Aggregation and Scale-Aware Regression for Monocular 3D Object Detection

Abstract
Monocular 3D object detection has received considerable attention for its simplicity and low cost. Existing methods typically follow conventional 2D detection paradigms, first locating object centers and then predicting 3D attributes via neighboring features. However, these approaches mainly focus on local information, which may limit the model’s global context awareness and result in missed detections, as the global context provides semantic and spatial dependencies essential for detecting small objects in cluttered or occluded environments. In addition, due to large variation in object scales across different scenes and depths, inaccurate receptive fields often lead to background noise and degraded feature representation. To address these issues, we introduce MonoASRH, a novel monocular 3D detection framework composed of Efficient Hybrid Feature Aggregation Module (EH-FAM) and Adaptive Scale-Aware 3D Regression Head (ASRH). Specifically, EH-FAM employs multi-head attention with a global receptive field to extract semantic features and leverages lightweight convolutional modules to efficiently aggregate visual features across different scales, enhancing small-scale object detection. The ASRH encodes 2D bounding box dimensions and then fuses scale features with the semantic features aggregated by EH-FAM through a scale-semantic feature fusion module. The scale-semantic feature fusion module guides ASRH in learning dynamic receptive field offsets, incorporating scale information into 3D position prediction for better scale-awareness. Extensive experiments on the KITTI and Waymo datasets demonstrate that MonoASRH achieves state-of-the-art performance. The code and model are released at https://github.com/WYFDUT/MonoASRH.
Index Terms:
3D object detection, monocular, scale-aware, scene understanding, autonomous driving.
I Introduction
In recent years, 3D object detection has emerged as a pivotal area of research, particularly driven by its critical role in autonomous systems, robotics, and augmented reality. The ability to accurately detect and locate objects in three-dimensional space is essential for applications that require a high degree of spatial awareness, especially for autonomous driving systems[1],[2].
Advancements in sensors such as LiDAR, radar, and stereo cameras have enhanced 3D object detection with point cloud data, radar signals, and depth maps. However, LiDAR’s high-resolution point clouds[3],[4],[5],[6] and stereo methods’ depth maps[7],[8] require additional calibration. CAD[9],[10] and multi-frame[11] approaches improve shape and movement tracking, but are also computationally intensive. Compared to the above methods, monocular vision systems offer a more cost-effective and deployable solution. However, they cannot accurately recover 3D information from 2D images without explicit depth perception.
Therefore, current advances in monocular 3D detection have focused on improving depth estimation accuracy. Methods like[12],[13],[14],[15],[16],[17] following the CenterNet[18] paradigm, use direct regression to estimate the 3D center depth of the target object.[9],[19],[20],[21],[22] incorporate geometric constraints into depth estimation to enhance target depth recovery.[23],[24],[25],[26],[27] effectively integrate geometric depth with regressed depth, improving the network’s generalization ability. However, these approaches focus on decoupling the regression of 2D and 3D attributes: predicting 2D bounding box parameters $\left({x,y,w,h}\right)$ and 3D object parameters $\left({x,y,z,w,h,l,yaw}\right)$ separately. This separation may neglect valuable spatial cues provided by the 2D predictions, such as the object’s size and position in the image, which can serve as geometric information for estimating depth and 3D structure. For instance, distant objects tend to produce smaller 2D bounding boxes, while nearby objects appear larger, indicating a potential correlation between 2D scale and 3D distance that is not fully exploited in decoupled frameworks. This issue is evident in the contrast between the attention heatmaps of the networks shown in Fig. 1. It can be seen that DEVIANT[22] and MonoLSS[17] struggle to focus on smaller, distant objects, particularly in the pedestrian category. Moreover, DEVIANT[22] tends to inadvertently take care of irrelevant background noise. Some previous works attempted to address this issue. Chen et al. proposed a shape-aware auxiliary training task[28]. [29] integrated deformable convolutions[30] to enhance the adaptability of the model to the features. Although these methods achieve dynamic receptive field adjustment to some extent, they have not explicitly considered object scale within a scene, nor dynamically adjusted the model’s attention based on different scales.
(a) DEVIANT
(b) MonoLSS
(c) MonoASRH
Figure 1: Visualization of attention heatmaps for different models: (a) DEVIANT[22], (b) MonoLSS[17], and (c) Our proposed MonoASRH. Previous methods struggle to capture distant and occluded objects due to fixed receptive fields. In contrast, MonoASRH dynamically adjusts attention across multiple scales, improving detection of various object categories, such as cars and pedestrians.
To address the aforementioned limitations, this paper proposes the novel Efficient Hybrid Feature Aggregation Module (EH-FAM) and Adaptive Scale-Aware 3D Regression Head (ASRH). Inspired by RT-DETR’s[31] success in 2D detection, EH-FAM integrates Vision Transformers with CNNs to efficiently aggregate visual features, significantly reducing model complexity compared to traditional methods like DLAUp[32]. The EH-FAM employs self-attention mechanisms for within-scale feature interaction at the highest level, while effectively propagating low-level feature information across all scales through a dedicated CNN architecture. Subsequently, ASRH encodes 2D bounding box dimensions to capture scale features, which are fused with semantic features extracted from the cross-scale feature through a specially designed scale-semantic feature fusion module. The fused features guide the learning of receptive field offsets, which are then applied to deformable convolutions, enhancing the model’s ability to adapt to varying object scales. Furthermore, ASRH utilizes a spatial variance-based attention mechanism to adaptively differentiate foreground objects from noise within semantic features. Additionally, we introduce the Selective Confidence-Guided Heatmap Loss to facilitate ASRH to prioritize high confidence detections and mitigate the impact of hard samples.
Our contributions can be summarized as follows:
  • We introduce a plug-and-play module called Efficient Hybrid Feature Aggregation Module, which is designed for efficient cross-scale feature aggregation.
  • We propose a novel Adaptive Scale-Aware 3D Regression Head, which dynamically adjusts the network’s receptive field according to object scale. By leveraging 2D predictions, ASRH facilitates 3D bounding box regression.
  • To enhance training stability, we introduce a Selective Confidence-Guided Heatmap Loss that emphasizes high-confidence detections while mitigating the impact of hard samples.
  • Extensive experiments on the KITTI 3D object detection benchmark[33] and Waymo Open dataset[34] show that our MonoASRH achieves better results compared to previous state-of-the-art methods.
II Related Works
II-A Monocular 3D Object Detection
In general, Monocular 3D object detection methods can be categorized into three main approaches: depth map-based, center-based, and transformer-based methods. Depth map-based approaches address spatial ambiguity by estimating depth maps using supervised [35],[36] or self-supervised networks[37], [38].[39], with techniques like pseudo-LiDAR conversions[40],[41] and multi-scale depth fusion[42],[43] improving geometric reasoning. However, their reliance on pre-trained depth networks and overlap between training and validation datasets[44] often lead to foreground depth inaccuracies.
Center-based methods extend anchor-free frameworks like CenterNet [18], linking 3D attributes to object centers. To mitigate depth errors, works like[14],[20],[22] integrate geometric constraints and hierarchical learning, while[24] leverages multiple depth attributes to generate various depth estimates. [28],[45] propose shape-aware schemes and plane-constrained 3D detection frameworks, respectively. MonoCD[27] introduces the concept of complementary depth. MonoLSS[17] uses Gumbel-Softmax probabilistic sampling[46] to distinguish between positive and negative samples in depth maps. Despite efficiency, these methods risk overlooking 2D-3D geometric consistency.
Transformer-based models leverage end-to-end architectures[47],[5] for global feature interaction, injecting global depth information into the Transformer to guide detection. MonoDETR[48] interacts with depth and visual features through a depth cross-attention layer in the decoder. To improve inference efficiency, MonoATT[49] introduces an adaptive token Transformer. However, their computational complexity, sensitivity to background noise, and slow convergence hinder real-time applicability.
In this work, we propose a center-based framework that integrates lightweight convolutional modules with multi-head attention for feature aggregation. To enhance 3D detection robustness, we introduce a latent space projection paradigm where 2D scale features are mapped to guide the 3D regression heads.
II-B Feature Aggregation Strategy
Most center-based monocular 3D detection frameworks[14],[20],[22],[27],[17] employ DLAUp[32] as the feature aggregation module. DLAUp’s hierarchical aggregation through pure convolution operations lacks sufficient global perception, and its large-kernel transposed convolution for upsampling also increases network parameters. Although MonoDETR[48] introduces an end-to-end Transformer architecture, traditional Transformer Encoder’s sensitivity to background noise and slow convergence hinder its performance.
Recent studies focus on hybrid CNN-Transformer architectures for feature aggregation. Works like U-shaped hybrid network by[50], cross-scale token attention by[51], dual-resolution self-attention by[52], and RT-DETR’s[31] efficient hybrid encoder demonstrate their effectiveness in modeling global dependencies while preserving local feature extraction. This hybrid architecture can effectively mitigate the aforementioned challenges in feature aggregation modules of monocular 3D detection methods.
Inspired by RT-DETR, our EH-FAM employs multi-head attention on highest-level features for long-range dependency capture. We further design a lightweight CNN-based cross-scale fusion structure that significantly enhances inference efficiency through re-parameterization[53] and convolutional decoupling[54] techniques.
II-C Scale-Aware Object Detection
Scale-aware techniques enhance detection models by incorporating scale information. SAFR-CNN[55] introduces scale-aware weighting to prioritize sub-networks specialized for input scales. SFDet[56] adapts this concept to face detection by associating multi-scale layers with appropriately sized anchors. Chen et al.[57] develop Scale-aware AutoAug for learning data augmentation strategies, while AdaZoom[58] proposes adaptive region zooming based on object scales. SMT[59] advances this field with Scale-Aware Modulation, integrating multi-head hybrid convolutions with lightweight aggregation for context-aware token refinement.
Currently, most methods focus on scale-awareness in 2D object detection, whereas there is relatively little work on monocular 3D detection. In fact, scale information is crucial for 3D detectors, since an object’s 2D size often has an implicit geometric projection relationship with its 3D position. Our ASRH module encodes 2D scale cues to generate dynamic receptive field offsets that guide the regression head toward scale-aware 3D localization.
III Proposed Methodology
In this section, we introduce our MonoASRH architecture. The overall framework, shown in Fig. 2, mainly consists of backbone, Efficient Hybrid Feature Aggregation Module, 2D regression head, and Adaptive Scale-Aware 3D Regression Head. Our pipeline mainly builds on the GUPNet [21]. Detailed implementation will be discussed later in this section.
Figure 2: Overview of our framework. The Efficient Hybrid Feature Aggregation Module (EH-FAM) efficiently aggregates multi-scale features. The Adaptive Scale-Aware 3D Regression Head (ASRH) fuses scale features with local semantics to guide the learning of 3D regression head.
III-A Overall Architecture
For a given RGB image $I\in{\mathbb{R}^{H\times W\times 3}}$, we extract multi-scale deep features using a pre-trained DLA-34 backbone[32]. These features, represented as $\left\{{{S_{1}},{S_{2}},{S_{3}},{S_{4}}}\right\}$, undergo further refinement and cross-scale fusion through the proposed EH-FAM. The resulting aggregated features are denoted as $F\in{\mathbb{R}^{\frac{H}{4}\times\frac{W}{4}\times C}}$. Similar to[21], our 2D detector is based on CenterNet[18]. We feed the deep features $F$ into three 2D detection heads to regress heatmaps $H\in{\mathbb{R}^{\frac{H}{4}\times\frac{W}{4}\times 3}}$, 2D offsets ${O_{2D}}\in{\mathbb{R}^{\frac{H}{4}\times\frac{W}{4}\times 2}}$, and 2D sizes ${S_{2D}}\in{\mathbb{R}^{\frac{H}{4}\times\frac{W}{4}\times 2}}$, enabling the prediction of 2D object centers and dimensions.
The deep features $F$ are also fed into ASRH for 3D attributes prediction. Specifically, RoI-Align extracts local semantic features ${F_{R}}\in{\mathbb{R}^{n\times d\times d\times C}}$ from $F$ based on 2D boxes, where $d\times d$ is the RoI-Align size and $n$ is the number of regions of interest. Subsequently, we denote the size of 2D bounding box corresponding to each region of interest as ${S_{R}}\in{\mathbb{R}^{n\times 2}}$, which is further encoded into scale features. Scale features and semantic features are then fused through the scale-semantic feature fusion module. Ultimately, ASRH outputs 3D bounding box sizes ${S_{3D}}\in{\mathbb{R}^{n\times 3}}$, 3D center offsets ${O_{3D}}\in{\mathbb{R}^{n\times 2}}$, yaw angle $\Theta\in{\mathbb{R}^{n\times 24}}$, direct depth $D\in{\mathbb{R}^{n\times d\times d}}$, and depth uncertainty ${D_{U}}\in{\mathbb{R}^{n\times d\times d}}$. Additionally, a depth attention map ${D_{M}}\in{\mathbb{R}^{n\times d\times d}}$ is employed to reduce the impact of irrelevant information, improving 3D detection accuracy.
III-B Efficient Hybrid Feature Aggregation Module
To efficiently aggregate features extracted by backbone at different scales into deep representations, we propose EH-FAM, a plug-and-play cross-scale feature aggregator. The “hybrid” module combines attention-based intra-scale interactions with CNN-based cross-scale fusion for improved performance and computational efficiency. The detailed implementation of EH-FAM is shown in Fig. 3.
Figure 3: Efficient Hybrid Feature Aggregation Module. Self-attention is applied only to the highest-level semantic feature map using an 8-head multi-head attention mechanism. Besides, all convolutional layers use the Mish activation function for improved feature representation.
1) Self-Attention Block: The EH-FAM receives four feature maps with different downsampling ratios as input. To ensure global extraction of rich semantic information about objects, we apply multi-head self-attention to the highest-level feature ${S_{4}}$. We use separate linear projections for queries, keys, and values:
${{Q_{i}}=\left({{X_{seq}}+{P_{e}}}\right)W_{i}^{\left(Q\right)},}&{{K_{i}}=\left({{X_{seq}}+{P_{e}}}\right)W_{i}^{\left(K\right)},}\end{array}$ (1)
${V_{i}}={X_{seq}}W_{i}^{\left(V\right)},$ (2)
where ${{{X_{seq}}}}\in{\mathbb{R}^{T\times C}}$ obtained by flattening ${S_{4}}$, $T=\frac{H}{{32}}\times\frac{W}{{32}}$ represents sequence length, ${{P_{e}}}$ is the positional embedding vector, and $W_{i}^{\left(Q\right)},W_{i}^{\left(K\right)},W_{i}^{\left(V\right)}$ are the query, key, and value transformation matrices for the $i$-th head, respectively. For each head, the attention output is a weighted sum of the values:
${Z_{i}}=\operatorname{Softmax}\left({\frac{{{Q_{i}}K_{i}^{T}}}{{\sqrt{d_{v}}}}}\right){V_{i}}.$ (3)
Finally, the concatenated output is passed through a linear layer to combine the multiple heads into a single output:
${Z_{out}}=[{Z_{1}},{Z_{2}},\ldots,{Z_{h}}]{W_{o}},$ (4)
where ${W_{o}}\in{\mathbb{R}^{\left({h\times{d_{v}}}\right)\times C}}$ is a learnable projection matrix. $h$ and $d_{v}$ denote the number of attention heads and the channel dimensions of ${{V_{i}}}$, respectively. Next, the output of the multi-head attention is passed through an FFN and then reshaped back to its original spatial dimensions, denoted as ${S_{out}}$.
2) Upsampling Block: To improve EH-FAM’s ability to perceive small-scale objects, we design an upsample block and fusion block composed of lightweight convolutional modules. These blocks enable cross-scale information interaction and enhance the extraction of local semantic features. In the upsample block, we replace DLAUp’s[32] large-kernel transposed convolutions with bilinear upsampling. To further mitigate information loss, we apply a 7×7 convolution prior to upsampling, which effectively refines feature maps. For efficiency, we decouple the 7$\times$7 convolution into successive 1$\times$7 and 7$\times$1 convolutions, following the approach in[54]. Additionally, since some categories, such as Pedestrian and Cyclist, have relatively large aspect ratios but small sizes, decoupled convolutions in the horizontal and vertical directions are adopted to capture these elongated scale features.
3) Fusion Block: In the fusion block, we employ the RepVGGplus block[53], which leverages re-parameterization techniques to convert a multi-branch ResNet-style architecture into a single-path VGG-like model during inference. This maintains the representational power of the model while accelerating inference. The final deep features are aggregated as follows:
$F={f}\left({{S_{out}},{S_{3}},{S_{2}},{S_{1}}}\right),$ (5)
where ${f}$ denotes the upsampling and cross-scale fusion operations.
III-C Adaptive Scale-Aware 3D Regression Head
As illustrated in Fig. 4, the proposed Adaptive Scale-Aware 3D Regression Head decomposes the 3D bounding box regression process into three stages. First, scale features are captured by encoding the 2D bounding box dimensions. Next, a scale-semantic fusion module combines these scale features with semantic features extracted from the region of interest, dynamically adjusting the 3D regression head’s receptive field. Given the limited image space occupied by foreground objects (e.g., the car class in KITTI covers just 11.42% of depth pixels), we also introduce an attention mask to ensure the 3D regression head focuses on relevant foreground regions. Finally, the 3D regression head outputs the 3D bounding box attributes.
Figure 4: The proposed Adaptive Scale-Aware 3D Regression Head. In Scale-Aware Offset, brown arrows indicate the direction of dynamic adjustment of the convolutional receptive field. In Attention Mask, varying shades of blue represent different weights at each position in the convolution kernel. Our network includes six parallel 3D regression heads of this structure to regress different 3D attributes.
1) Scale Encoder: The Scale Encoder uses MLP to transform 2D geometric properties into a high-dimensional representation. By learning this transformation, the Scale Encoder effectively encodes scale information. For each region of interest, the corresponding 2D bounding box size ${S_{R}}\in{\mathbb{R}^{n\times 2}}$ is mapped to a high-dimensional feature space:
${G^{\prime}_{o}}=\operatorname{MLP}\left({{S_{R}}}\right),$ (6)
where ${G^{\prime}_{o}}\in{\mathbb{R}^{n\times 32}}$. Simultaneously, we apply a series of convolutional layers to further refine the feature maps ${F^{\prime}_{R}}$ within the region of interest, obtaining the local semantic features ${F_{R}}$:
${F_{R}}=\operatorname{Conv}\left({\operatorname{ReLU}\left({\operatorname{Conv}\left({{F^{\prime}_{R}}}\right)}\right)}\right),$ (7)
where ${F_{R}}\in{\mathbb{R}^{n\times d\times d\times C}}$, $d\times d$ is the RoI-Align size. Offset $G^{\prime}_{o}$ is then reshaped to match the spatial dimensions of the ${F_{R}}$, denoted as ${G_{o}}$.
2) Scale-Semantic Feature Fusion Module: This module comprises two key components: offset feature generation and attention mask generation. For offset generation, the scale feature ${G_{o}}$ and local semantic features ${F_{R}}$ are concatenated. The combined features are processed through a stacked convolutional layer consisting of 3$\times$3 and 1$\times$1 convolutions, producing a scale-aware offset ${\Delta_{S}}$, which is used to dynamically adjust the receptive field in subsequent deformable convolutions.
${\Delta_{S}}=\operatorname{Conv}\left({\operatorname{ReLU}\left({\operatorname{Conv}\left({\left[{{G_{o}},{F_{R}}}\right]}\right)}\right)}\right),$ (8)
where ${\Delta_{S}}\in{\mathbb{R}^{n\times d\times d\times 18}}$, and 18 represents the offset of each spatial position in the deformable convolution kernel.
In the attention mask generation component, inspired by[60], we design an attention mechanism based on spatial variance. This mechanism emphasizes regions with significant deviations from the mean, which are often foreground regions of interests. First, the mean of the local semantic features ${F_{R}}$ are computed across spatial dimensions:
$\mu=\frac{1}{{d\times d}}\sum\limits_{i=1}^{d}{\sum\limits_{j=1}^{d}{F_{R}^{n,i,j,c}}}.$ (9)
Then, we calculate the squared difference of each pixel relative to the mean across the spatial dimensions of the feature map:
$V={\left({F_{R}^{n,i,j,c}-\mu}\right)^{2}}.$ (10)
Using the squared difference obtained from the Eqn. 10, the attention weights can be calculated as:
$Y=\sigma\left({\frac{V}{{{\frac{1}{{d\times d-1}}\sum\limits_{i=1}^{d}{\sum\limits_{j=1}^{d}{{V_{i,j}}+\lambda}}}}}+0.5}\right),$ (11)
where $\lambda$ is a small positive constant. $\sigma$ is a projection function implemented by convolution followed by sigmoid. Normalizing by the variance provides a measure of how relatively important a specific deviation is compared to the overall distribution of deviations.
Finally, a convolutional layer with a sigmoid activation function is applied to generate the mask:
$M=\sigma\left({{F_{R}}\odot Y}\right),$ (12)
where $M\in{\mathbb{R}^{n\times d\times d\times 9}}$, and $\odot$ denotes element-wise multiplication.
3) 3D Regression Head: To achieve scale-aware dynamic receptive field adjustment, the first layer of the 3D regression head utilizes deformable convolutions[30]. As shown in Eqn. 13, the offset ${\Delta_{S}}$ and attention mask $M$ generated by the Scale-Semantic Feature Fusion Module are applied to this layer, allowing the model to better detect and localize objects of varying scales:
${F_{de}}=\sum\limits_{{P_{n}}\in\Omega}{{M_{{P_{n}}}}\cdot{W_{{P_{n}}}}\cdot{F_{R}}\left({\Delta_{S}^{{P_{n}}}+{P_{n}}}\right)},$ (13)
where $\Omega$ represents the local neighborhood for a 3$\times$3 kernel, ${{P_{n}}}$ enumerates the locations in $\Omega$, ${\Delta_{S}^{{P_{n}}}}$ is the learned offset for each position in the convolution kernel, ${{W_{{P_{n}}}}}$ is the weight of the convolution kernel at position ${{P_{n}}}$, and ${{M_{{P_{n}}}}}$ controls the contribution of the corresponding kernel position.
Next, we normalize the feature map ${F_{de}}$ using the Attentive Normalization layer[61], which is a light weight module integrating feature normalization and channel-wise feature attention:
${F_{n}}=\operatorname{AN}\left({{F_{de}}}\right).$ (14)
Due to AN’s mixture modeling approach to affine transformation, which recalibrates features after standardization, it is employed in the regression head to facilitate the learning of more expressive latent feature representations. Finally, the feature map is processed through a LeakyReLU activation function and a 1$\times$1 convolution for channel mapping, ultimately regressing the 3D bounding box attributes.
III-D Loss Function
1) Selective Confidence-Guided Heatmap Loss: Easy samples, which are typically fully visible and closer to the camera, exhibit less scale complexity. As a result, our scale-aware mechanism struggles to provide additional focus on these targets. To address this issue, we propose the Selective Confidence-Guided Heatmap Loss, which encourages the model to focus more on high-confidence samples, given that the network tends to assign higher confidence scores to easy samples (as shown in Fig. 5). Specifically, this loss function uses an improved version of Focal Loss[62] as the primary loss and introduces a Selective Confidence-Guided (SCG) Loss as an auxiliary loss (Eqn. 15) to enhance training.
Figure 5: We visualize the confidence score distribution of MonoASRH across three difficulty levels in the KITTI dataset, considering only samples with a 3D IoU greater than 0.5. The model tends to assign higher confidence scores to easy samples, while showing lower confidence for hard samples.
${\mathcal{L}_{SCG}}=-\frac{1}{B}\sum\limits_{i=1}^{B}{\sum\limits_{j=1}^{K}\left\{\begin{aligned} &\log\left({{t_{ij}}}\right),&{t_{ij}}>T\\ &0,&{\text{otherwise}}\\ \end{aligned}\right.},$ (15)
where $B$ is the batch size, $K$ is the maximum number of detectable objects per image, ${t_{ij}}$ are the top $K$ points with the highest confidence in the predicted heatmap, and $T$ is the confidence threshold.
The focal loss introduces a dynamic weighting for heatmap regression. The SCG Loss encourages the ASRH to prioritize high-confidence samples, which is critical for accurate detection of closer objects. Meanwhile, optimizing for high-confidence objects can to some extent mitigate the overfitting of noise in low-confidence samples. The final heatmap regression loss is a weighted combination of these two losses:
${\mathcal{L}_{Heatmap}}={\mathcal{L}_{focal}}+\lambda{\mathcal{L}_{SCG}}.$ (16)
In the experiment, $T$ was taken as 0.9 and $\lambda$ as 0.01.
2) Total Loss: The overall network loss $\mathcal{L}$ consists of two components: 2D object loss ${\mathcal{L}_{2D}}$ and 3D object loss ${\mathcal{L}_{3D}}$. Mathematically, the composite loss function is expressed as:
$\mathcal{L}={\mathcal{L}_{2D}}+{\mathcal{L}_{3D}}.$ (17)
The 2D object loss follows the design of CenterNet[18]. It includes the Selective Confidence-Guided Heatmap Loss ${\mathcal{L}_{Heatmap}}$, 2D center offset loss ${\mathcal{L}_{{O_{2D}}}}$ and 2D size loss ${\mathcal{L}_{{S_{2D}}}}$:
${\mathcal{L}_{2D}}={\mathcal{L}_{Heatmap}}+{\mathcal{L}_{{O_{2D}}}}+{\mathcal{L}_{{S_{2D}}}},$ (18)
where ${\mathcal{L}_{{O_{2D}}}}$ and ${\mathcal{L}_{{S_{2D}}}}$ use standard L1 loss.
Following the MonoDLE paradigm[14], the 3D object loss trains the network to predict key 3D attributes. It includes the 3D center offset loss ${\mathcal{L}_{{O_{3D}}}}$ and 3D size loss ${\mathcal{L}_{{S_{3D}}}}$ with L1 loss, the orientation loss ${\mathcal{L}_{\theta}}$ using multi-bin loss[63], and the depth loss ${\mathcal{L}_{D}}$ supervised by Laplacian aleatoric uncertainty loss[64]:
${\mathcal{L}_{3D}}={\mathcal{L}_{{O_{3D}}}}+{\mathcal{L}_{{S_{3D}}}}+{\mathcal{L}_{\theta}}+{\mathcal{L}_{D}}.$ (19)
IV Experiments
IV-A Setup
1) Dataset: To evaluate the performance of our proposed method, we conducted experiments on the KITTI 3D object detection benchmark[33] and Waymo Open dataset[34]. The KITTI dataset consists of 7,481 training images and 7,518 testing images. It includes three primary object classes: Car, Pedestrian, and Cyclist, with varying levels of difficulty (Easy, Moderate, and Hard) based on factors such as object size, occlusion, and truncation. Following the previous studies[65], the training images are further split into a training set of 3,712 images and a validation set of 3,769 images.
The Waymo Open Dataset[34] is a pivotal resource for autonomous driving scene understanding, featuring 798 training and 202 validation sequences, yielding approximately 160,000 and 40,000 samples, respectively. For our experiments, we adopt the approach from[65], generating 52,386 training images and 39,848 validation images by sampling every third frame from the training sequences. This dataset captures diverse real-world driving scenarios and categorizes objects into LEVEL_1 and LEVEL_2 based on LiDAR point density.
2) Evaluation metrics: In KITTI, model’s performance is typically measured using average precision in 3D space and birds-eye view ($A{P_{3D}}$ and $A{P_{BEV}}$) at 40 recall positions. The AP is calculated based on Intersection over Union between the predicted 3D bounding boxes and the ground truth, with specific IoU thresholds for each object class – 0.7 for cars and 0.5 for pedestrians and cyclists. Following the official protocol[33], we use $A{P_{3D}}$ and $A{P_{BEV}}$ on Moderate category as main metrics. In Waymo, evaluations use two IoU thresholds (0.5 and 0.7) across four distance ranges: Overall, 0 - 30m, 30 - 50m, and 50m - $\infty$.
3) Data augmentation: To enhance the generalization ability of our model, we incorporated several data augmentation techniques commonly used in object detection. During training, we applied random horizontal flipping and random cropping with a probability of 0.5, random scaling with a probability of 0.4, and random shifting with a probability of 0.1. Additionally, we utilized the MixUp3D technique from MonoLSS[17] as an extra data augmentation strategy to further improve the model’s robustness against occlusions.
4) Implementation details: Our proposed MonoASRH was trained on two RTX 3090 GPUs with batch size of 16. For the EH-FAM, we set the channel dimension ${d_{v}}$ of ${V_{i}}$ to 128 in the multi-head self-attention. Following[21], the RoI-Align size $d\times d$ is set to $7\times 7$. For KITTI, we trained the model for 450 epochs using the Adam optimizer[66], with a weight decay of $51e-$ 5. The initial learning rate was set to $31e-$ 3 and decayed by a factor of 10 at the 250th and 370th epochs. For Waymo Open dataset, we trained the model for only 80 epochs with a batch size of 32, while keeping all other settings consistent with those used for KITTI. During testing, the maximum number of detectable objects per image, $K$, was set to 50.
IV-B Main Results
We compared our proposed MonoASRH with state-of-the-art monocular 3D object detection methods on the KITTI test set for the car category. As shown in Table I, our method outperformed the CAD model-based DCD[10], even without additional data. Compared to other methods that do not use extra data, MonoASRH achieved improvements in $A{P_{3D|R40}}$ across all three difficulty levels by 0.65%, 1.00%, and 1.40%. Experimental results validate the effectiveness of our method in detecting objects in the Car category.
TABLE I: Comparisons for the Car Category on the KITTI test set at $\operatorname{IoU}=0.7$. The Best and Second Best Results are Bolded and Underlined, Respectively. The Improvements are Shown in Purple Fonts. ‘’ denote Reproduced Results.
Methods Reference Extra Data Test, $A{P_{3D|R40}}$(%) Test, $A{P_{BEV|R40}}$(%) Runtime (ms)
Easy Mod. Hard Easy Mod. Hard
D4LCN[39] CVPR 2020 Depth 16.65 11.72 9.51 22.51 16.02 12.55 -
Kinematic3D[11] ECCV 2020 Multi-frames 19.07 12.72 9.17 26.99 17.52 13.10 120
MonoRUn[3] CVPR 2021 LiDAR 19.65 12.30 10.58 27.94 17.34 15.24 70
CaDDN[4] CVPR 2021 LiDAR 19.17 13.41 11.46 27.94 18.91 17.19 630
AutoShape[9] ICCV 2021 CAD 22.47 14.17 11.36 30.66 20.08 15.59 -
MonoDTR[5] CVPR 2022 LiDAR 21.99 15.39 12.73 28.59 20.38 17.14 37
DCD[10] ECCV 2022 CAD 23.81 15.90 13.21 32.55 21.50 18.25 -
SMOKE[13] CVPRW 2020 None 14.03 9.76 7.84 20.83 14.49 12.75 30
MonoDLE[14] CVPR 2021 None 17.23 12.26 10.29 24.79 18.89 16.00 40
GUPNet[21] ICCV 2021 None 20.11 14.20 11.77 - - - 34
MonoGround[25] CVPR 2022 None 21.37 14.36 12.62 30.07 20.47 17.74 30
DEVIANT[22] ECCV 2022 None 21.88 14.46 11.89 29.65 20.44 17.43 40
MonoCon[16] AAAI 2022 None 22.50 16.46 13.95 31.12 22.10 19.00 26
MonoDDE[24] CVPR 2022 None 24.93 17.14 15.10 33.58 23.46 20.37 40
PDR[67] TCSVT 2023 None 23.69 16.14 13.78 31.76 21.74 18.79 29
DVDET[68] RAL 2023 None 23.19 15.44 13.07 32.05 22.15 19.32 -
MonoDETR[48] ICCV 2023 None 25.00 16.47 13.58 33.60 22.11 18.60 38
MonoLSS[17] 3DV 2024 None 24.91 18.17 15.52 34.90 25.65 22.45 35
MonoCD[27] CVPR 2024 None 25.53 16.59 14.53 33.41 22.81 19.57 36
FD3D[69] AAAI 2024 None 25.38 17.12 14.50 34.20 23.72 20.76 40
MonoASRH(Ours) - None 26.18 19.17 16.92 34.84 25.89 22.48 32
Improvement - v.s. second-best +0.65 +1.00 +1.40 -0.06 +0.24 +0.03 -
TABLE II: Comparisons for the Pedestrian and Cyclist Category on the KITTI test set at $\operatorname{IoU}=0.5$. The Best and Second Best Results are Bolded and Underlined, Respectively.
Methods Extra Data Test, $A{P_{3D|R40}}$(%)
Pedestrian Cyclist
Easy Mod. Hard Easy Mod. Hard
CaDDN[4] LiDAR 12.87 8.14 6.76 7.00 3.41 3.30
MonoDTR[5] LiDAR 15.33 10.18 8.61 5.05 3.27 3.19
DCD[10] CAD 10.37 6.73 6.28 4.27 2.74 2.41
DEVIANT[22] None 13.43 8.65 7.69 5.05 3.13 2.59
MonoCon[16] None 13.10 8.41 6.94 2.80 1.92 1.55
MonoDDE[24] None 11.13 7.32 6.67 5.94 3.78 3.33
PDR[67] None 11.61 7.72 6.40 2.72 1.57 1.50
MonoDETR[48] None 12.54 7.89 6.65 7.33 4.18 2.92
MonoASRH(Ours) None 16.90 11.24 9.64 6.49 3.94 3.51
TABLE III: Performance of the Car Category on the KITTI validation set at $\operatorname{IoU}=0.7$. The Best and Second Best Results are Bolded and Underlined, Respectively.
Methods Val, $A{P_{3D|R40}}$(%) Val, $A{P_{BEV|R40}}$(%)
Easy Mod. Hard Easy Mod. Hard
GUPNet[21] 22.76 16.46 13.72 31.07 22.94 19.75
DEVIANT[22] 24.63 16.54 14.52 32.60 23.04 19.99
MonoDDE[24] 26.66 19.75 16.72 35.51 26.48 23.07
PDR[67] 27.65 19.44 16.24 35.59 25.72 21.35
MonoDETR[48] 28.84 20.61 16.38 - - -
MonoCD[27] 26.45 19.37 16.38 34.60 24.96 21.51
MonoLSS[17] 25.91 18.29 15.94 34.70 25.36 21.84
FD3D[69] 28.22 20.23 17.04 36.98 26.77 23.16
MonoASRH(Ours) 28.35 20.75 17.56 37.53 27.26 24.16
The results for the Pedestrian and Cyclist categories in the KITTI test set are presented in Table II. For Pedestrians, MonoASRH achieved the best performance across all difficulty levels, outperforming the LiDAR-based MonoDTR method[5]. In the Cyclist category, MonoASRH ranked third at the Easy level but still showed competitive performance, with a slight drop at the Moderate level compared to the Transformer-based MonoDETR[48]. However, it achieved the best performance at the Hard level with a 3.51% $A{P_{3D|R40}}$. These results highlight the challenges of detecting smaller, non-rigid objects like Pedestrians and Cyclists. Nonetheless, MonoASRH’s scale-aware dynamic receptive field adjustment capability enabled strong performance in these categories, validating the model’s generalization and scalability. Additionally, the decoupled convolutional layers in EH-FAM are better equipped to capture the features of these categories with large aspect ratios.
Figure 6: Qualitative results produced by the proposed MonoASRH on the KITTI test set. The detected 3D bounding boxes for Car, Pedestrian and Cyclist are shown in purple, orange and green, respectively.
TABLE IV: Comparisons for the Vehicle Category on the Waymo Open validation set. We Use $A{P_{3D}}$ (LEVEL_1 and LEVEL_2, IoU $>$ 0.5 and IoU $>$ 0.7) according to Three Object Distance Intervals. ‘’ denote Reproduced Results.
Methods Reference LEVEL_1(IoU $>$ 0.5) LEVEL_2(IoU $>$ 0.5) LEVEL_1(IoU $>$ 0.7) LEVEL_2(IoU $>$ 0.7)
Overall 0-30m 30-50m 50m-Inf Overall 0-30m 30-50m 50m-Inf Overall 0-30m 30-50m 50m-Inf Overall 0-30m 30-50m 50m-Inf
With extra data:
PatchNet[43] ECCV 2020 2.92 10.03 1.09 0.23 2.42 10.01 1.07 0.22 0.39 1.67 0.13 0.03 0.38 1.67 0.13 0.03
PCT[70] NIPS 2021 4.20 14.70 1.78 0.39 4.03 14.67 1.74 0.36 0.89 3.18 0.27 0.07 0.66 3.18 0.27 0.07
NF-DVT[71] JAS 2024 11.32 23.35 5.76 0.98 11.24 23.20 5.68 0.92 2.76 5.89 0.99 0.11 2.64 5.85 0.96 0.10
Without extra data:
GUPNet[21] ICCV 2021 10.02 24.78 4.84 0.22 9.39 24.69 4.67 0.19 2.28 6.15 0.81 0.03 2.14 6.13 0.78 0.02
MonoJSG[72] CVPR 2022 5.65 20.86 3.91 0.97 5.34 20.79 3.79 0.85 0.97 4.65 0.55 0.10 0.91 4.64 0.55 0.09
DEVIANT[22] ECCV 2022 10.98 26.85 5.13 0.18 10.29 26.75 4.95 0.16 2.69 6.95 0.99 0.02 2.52 6.93 0.95 0.02
MonoRCNN++[73] WACV 2023 11.37 27.95 4.07 0.42 10.79 27.88 3.98 0.39 4.28 9.84 0.91 0.09 4.05 9.81 0.89 0.08
MonoUNI[74] NIPS 2023 10.98 26.63 4.04 0.57 10.38 26.57 3.95 0.53 3.20 8.61 0.87 0.13 3.04 8.59 0.85 0.12
MonoCD[27] CVPR 2024 11.62 27.21 5.24 0.83 11.14 27.14 5.12 0.75 3.85 8.43 1.00 0.13 3.50 7.53 0.87 0.11
MonoASRH(Ours) - 12.35 29.25 6.16 1.10 11.58 29.14 5.95 0.96 3.36 9.49 1.01 0.13 3.15 9.46 0.97 0.11
As shown in Table III, MonoASRH was also evaluated on the KITTI validation set. For the Car category at the 0.7 IoU threshold and Moderate difficulty, MonoASRH achieved state-of-the-art performance with $A{P_{3D|R40}}$ and $A{P_{BEV|R40}}$ scores of 20.75% and 27.26%, respectively. Some qualitative results on the KITTI test set are shown in Fig. 6. Benefiting from the globally-aware self-attention mechanism in EH-FAM, the model effectively captures long-range dependencies for smaller objects such as distant cars.
Table IV compares the $AP_{3D}$ metrics for the “Vehicle” category across various methods on the Waymo Open Val Set. In the 30-50m range, our model achieved state-of-the-art performance. Compared to NF-DVT[71], MonoASRH improved the $AP_{3D}$ by 0.4% and 0.02% at IoU thresholds of 0.5 and 0.7, respectively, within the 30-50m range under the LEVEL_1 setting. Additionally, under the LEVEL_2 setting, the $AP_{3D}$ metrics improved by 0.27% and 0.01% at different IoU thresholds within the 30-50m range. These results further validate the effectiveness of MonoASRH in detecting distant and small objects.
IV-C Ablation Studies
In this section, we evaluate the impact of each component of the proposed framework on the network’s performance. Ablation studies were conducted on the KITTI validation set using the Car category, with $A{P_{3D|R40}}$ and $A{P_{BEV|R40}}$ as metrics. Our baseline model is based on the GUPNet[21] and incorporates the Learnable Sample Selection from MonoLSS[17] for initial training. We compare the effects of different components, including the Adaptive Scale-Aware 3D Regression Head, Efficient Hybrid Feature Aggregation Module, and Selective Confidence-Guided Heatmap Loss. As shown in Fig. 7, we provide detailed visual comparisons of various approaches when the model is not equipped with a specific module. The results in Table V and Fig. 7 demonstrate the contributions of each module to the overall performance of our method.
Figure 7: Qualitative comparison results of ablation study on different components of our overall framework in the BEV view. Predicted and ground truth bounding boxes are shown in red and green, respectively.
Figure 8: Qualitative comparison results of ablation study on ASRH in the BEV view. Predicted and ground truth bounding boxes are shown in red and green, respectively.
(a) MonoLSS
(b) MonoDETR
(c) MonoASRH
Figure 9: Qualitative comparison results of 3D object detection on the KITTI val set: (a) MonoLSS[17], (b) MonoDETR[48], and (c) Ours. In each group of images, the first row shows the camera view, and the second row shows the BEV view. In the camera view, predicted and ground truth 3D bounding boxes are shown in purple and green, respectively. In the BEV view, their projections are represented in red and green.
TABLE V: Ablation Study on Different Components of our Overall Framework on KITTI val set for Car Category. ‘$A$’, ‘$E$’ and ‘${\mathcal{L}_{H}}$’ denote Adaptive Scale-Aware 3D Regression Head, Efficient Hybrid Feature Aggregation Module, and Selective Confidence-Guided Heatmap Loss, Respectively.
        $A$        $E$        ${\mathcal{L}_{H}}$        Val, $A{P_{3D|R40}}$(%)
               Easy        Mod.        Hard
        -        -        -        25.30        18.20        15.80
               -        -        26.12        20.32        17.34
        -        ✓        -        26.67        19.39        16.11
        -        -        ✓        25.87        18.23        15.76
               ✓        -        27.98        20.73        17.45
               ✓        ✓        28.35        20.75        17.56
1) Adaptive Scale-Aware 3D Regression Head: In Table VI, we further explore the contribution of each component within the Adaptive Scale-Aware 3D Regression Head to model’s performance. Specifically, we evaluated ASRH under three settings: “w/o Scale-Aware Offset”, where receptive field offsets are derived solely from semantic features without object scale information; “w/o Adaptive Attention Mask”, where the attention mask is randomly initialized; “w/o Attentive Normalization”, where the AN layer[61] is not applied to normalize feature maps ${F_{de}}$. The experimental results show that incorporating scale information is crucial in ASRH, enabling the model to perform dynamic scale-aware adjustments of the receptive field. This significantly improves detection, particularly for small and occluded objects at Moderate and Hard difficulty levels. We also provide detailed visual comparisons in Fig. 8, which further demonstrate the role of each component in ASRH.
TABLE VI: Ablation Study of ASRH on KITTI val set for Car Category.
       Architecture       Val, $A{P_{3D|R40}}$(%)
             Easy       Mod.       Hard
       Baseline + ASRH       26.12       20.32       17.34
       w/o Scale-Aware Offset       25.72       18.91       16.35
       w/o Adaptive Attention Mask       25.63       19.86       17.00
       w/o Attentive Normalization       25.97       20.15       17.16
Additionally, we compared ASRH with the standard deformable convolution[30]. We inserted deformable convolution into the first layer of each 3D regression head in the baseline model and conducted comparison experiments with ASRH. As shown in Table VII, ASRH consistently outperformed standard deformable convolutions in monocular 3D detection tasks.
TABLE VII: Comparison Experiment between ASRH and Deformable Convolution on KITTI val set for Car Category. ‘$\operatorname{Deform}.$’ Denotes Deformable Convolution.
Methods Val, $A{P_{3D|R40}}$(%) Val, $A{P_{BEV|R40}}$(%)
Easy Mod. Hard Easy Mod. Hard
Baseline 25.30 18.20 15.80 34.41 25.50 21.86
+ Deform. 25.71 18.85 16.88 34.55 25.80 22.27
+ ASRH 26.12 20.32 17.34 35.26 27.36 23.87
2) Efficient Hybrid Feature Aggregation Module: In order to validate EH-FAM’s generalizability, we replaced the original feature aggregation layers in three baseline models (MonoDLE[14], GUPNet[21], and MonoLSS[17]) with our EH-FAM:
  • MonoDLE uses a stacked hourglass network for multi-scale feature fusion. We replaced its final feature aggregation block with EH-FAM while retaining the backbone and detection heads.
  • GUPNet employs DLAUp[32], a hierarchical aggregation module. We substituted DLAUp entirely with EH-FAM, preserving the backbone and 3D regression head.
  • MonoLSS also adopts the DLAUp structure. We replaced it with EH-FAM, keeping the Learnable Sample Selection module unchanged.
No additional components were added or removed beyond this substitution. The baseline models’ training protocols, loss functions, and hyperparameters remained identical to ensure fair comparison.
The results in Table VIII show that EH-FAM led to performance improvements and parameter reduction across all three baselines. Specifically, the $A{P_{3D|R40}}$ scores for MonoDLE, GUPNet, and MonoLSS on the Moderate level increased by 0.75%, 0.05%, and 0.91%, respectively, while their parameter counts decreased by 8.76%, 8.74%, and 8.48%. Additionally, the computational cost was reduced by 5.71%, 7.25%, and 7.17%, respectively. These results confirm that EH-FAM can be widely applied to various baselines, significantly enhancing both performance and computational efficiency.
TABLE VIII: Ablation Study of EH-FAM with Different Baselines on KITTI val set for Car Category. $\oplus$’ Indicates that the Feature Aggregation Module of the Original Baseline Model is Replaced with EH-FAM.
Methods Val, $A{P_{3D|R40}}$(%) Params(M)$\downarrow$ FLOPs(G)$\downarrow$
Easy Mod. Hard
MonoDLE[14] 17.45 13.66 11.68 19.86 158.18
MonoDLE $\oplus$ EH-FAM 19.89 14.41 12.35 18.12 149.14
Improvement +2.44 +0.75 +0.67 -8.76% -5.71%
GUPNet[21] 22.76 16.46 13.72 19.90 124.57
GUPNet $\oplus$ EH-FAM 22.90 16.51 13.92 18.16 115.53
Improvement +0.14 +0.05 +0.20 -8.74% -7.25%
MonoLSS[17] 25.91 18.29 15.94 20.51 126.13
MonoLSS $\oplus$ EH-FAM 26.62 19.30 16.09 18.77 117.09
Improvement +0.71 +1.01 +0.15 -8.48% -7.17%
3) Selective Confidence-Guided Heatmap Loss: The Selective Confidence-Guided Heatmap Loss guides the network to focus more on high-confidence samples. As shown in Table V, when used alone, this loss achieves the largest improvement in the baseline model’s $A{P_{3D|R40}}$ for the Car category at the Easy level, reaching 25.87%. When using the other two modules, adding the loss further improves the $A{P_{3D|R40}}$ by 0.37%. This demonstrates that the loss function is particularly beneficial for detecting Easy samples in the KITTI dataset. The likely reason is that the loss encourages the model to rely more on high-confidence detections, and Easy samples are generally easier to detect, leading to higher confidence scores.
IV-D Visualization
In Fig. 9, we compare the proposed MonoASRH with the Transformer-based MonoDETR[48] and center-based MonoLSS[17]. Each group of images shows the camera view and BEV view. Compared to the other methods, MonoASRH produces higher-quality 3D bounding boxes across various scenes. Notably, it excels at detecting and localizing small distant objects, such as the white vehicle on the right in the third group of images, demonstrating the benefit of incorporating scale information. MonoASRH also provides more accurate results for fully visible objects. Additionally, it more effectively detects partially occluded objects, as shown on the left in the second group of images.
Figure 10: Qualitative results of failure cases on the KITTI val set. Predicted and ground truth 3D bounding boxes are shown in purple and green, respectively. The pentagrams represent objects that were not detected by the MonoASRH but were annotated in the ground truth.
V Limitation
While MonoASRH achieves state-of-the-art performance, it has several limitations. As Fig. 10 illustrates, MonoASRH may underperform when detecting heavily truncated objects (e.g., objects partially outside the image frame). We hypothesize that the aforementioned limitations may arise from the ASRH module’s reliance on 2D scale information extraction. Truncation often results in incomplete 2D bounding boxes, leading to inaccurate scale feature extraction by ASRH. This issue is particularly pronounced in crowded urban scenarios where truncation is common. Additionally, ASRH’s performance depends mainly on the accuracy of initial 2D object center predictions. Truncated objects often cause shifts in predicted center points, and errors in 2D detection (e.g., missed or mislocalized centers) propagate to the 3D regression stage, degrading overall performance.
Furthermore, while MonoASRH focuses on monocular 3D object detection, its potential in multi-view 3D detection tasks remains unexplored. By effectively bridging 2D and 3D detection, MonoASRH may demonstrate promising adaptability for multi-view 3D detection scenarios. The above limitations underscore future research directions, such as integrating uncertainty estimation for truncated objects, designing robust 2D-3D joint training strategies, and exploring extensions to multi-view 3D detection tasks.
VI Conclusion
We propose MonoASRH, which introduces Efficient Hybrid Feature Aggregation Module (EH-FAM) and Scale-Aware 3D Regression Head (ASRH) for monocular 3D detection. Our plug-and-play EH-FAM employs a hybrid architecture to efficiently aggregate features across different scales, endowing the model with richer scale-aware semantic information. ASRH dynamically adjusts the network’s receptive field according to the scale of objects, thereby enhancing 3D detection accuracy. Experimental results on the KITTI and Waymo benchmark demonstrate that MonoASRH achieves state-of-the-art performance. However, MonoASRH may underperform when detecting truncated objects, as these objects often lead to inaccuracies in the acquisition of scale features by ASRH. Future research will focus on overcoming this challenge.